What will be the nominal rate of return on a perpetual preferred stock with a $100 par value, a stated dividend of 10% of par, and a current market price of (a) $68, (b) $86, (c) $108, and (d) $137? Round your answers to two decimal places.
Answer #1

Annual dividend=100*10%=10

Hence nominal rate=annual dividend/current price

1.nominal rate=10/68=14.71%

2. Nominal rate=10/86=11.63%

3. Nominal rate=10/108=9.26%

4. Nominal rate=10/137=7.30%